# Context-Picker: Dynamic context selection using multi-stage reinforcement learning

### v1

@@ 4.2 Main Results
Method LoCoMo MultiFieldQA HotpotQA 2WikiMQA MuSiQue
LLM (Qwen3-8B, no retrieval) 0.566 0.833 0.661 0.389 0.280
RAG (Qwen3-8B)
0.622
(TopK=100)
0.857
(TopK=5)
0.857
(TopK=10)
0.597
(TopK=5)
0.700
(TopK=10)
0.525
(TopK=5)
0.560
(TopK=10)
0.340
(TopK=5)
0.390
(TopK=10)
Context-Picker, Stage 1 (Qwen3-8B) 0.681 0.873 0.741 0.621 0.476
Context-Picker, Stage 2 (Qwen3-8B) 0.706 0.825 0.747 0.702 0.522
Table 1: Main results on knowledge-intensive QA benchmarks. We report Judge Acc (higher is better). Best per column is in bold.
Comparison with LLM-only and RAG baselines.
Across all datasets, both stages of Context-Picker substantially outperform the non-retrieval LLM baseline, confirming that external evidence is crucial for long-context and multi-hop QA and that simply relying on parametric knowledge is insufficient in these settings. Even the recall-oriented Stage I, which tolerates some redundancy, already yields sizable gains over the plain LLM.
When compared under comparable evidence budgets, Context-Picker also brings consistent improvements over the vanilla RAG pipeline on most datasets. On LoCoMo, HotpotQA, 2WikiMQA, and MuSiQue, Stage 2 delivers the best overall performance, exceeding the strongest RAG baseline by +4+41818 points in Judge Acc. On MultiFieldQA, the recall-oriented Stage 1 slightly surpasses RAG (0.873 vs. 0.857), while Stage 2 trades a small drop in accuracy (0.825) for more compact inputs. These results suggest that, beyond a strong retriever, adaptively deciding which passages to keep and how many to include per query is beneficial: Context-Picker improves answer quality without simply increasing the number of passages and often reduces prompt overhead.
@@
Reinforcement learning on discrete text selection is often characterized by instability. However, thanks to our dense reward supervision mined via LOO and the GRPO algorithm, Context-Picker demonstrates robust training dynamics. As illustrated in Figure 3, the reward curves for both the Recall-Oriented Stage I (Left) and Precision-Oriented Stage II (Right) show steady convergence. The minimal gap between training and validation performance further validates the generalization capability of our offline evidence mining strategy.
@@ 4.3 Ablation Studies
In the full model, the picker outputs both a short natural-language rationale and a set of selected IDs. Removing the rationale branch (“w/o rationale”) leads to a $6.5$-point drop in Judge Acc and noticeably higher variance across runs. We hypothesize that requiring the model to verbalize why certain passages are selected acts as a structural regularizer: it encourages more stable reasoning over evidence interactions and reduces the tendency to over-select loosely related passages.
@@
Finally, we examine a variant trained only with the Stage II objective (“w/o Stage 1”), i.e., directly optimizing the refinement-oriented reward from scratch. This leads to a substantial degradation to $56.5\%$ Judge Acc, $14.1$ points below the full two-stage Context-Picker. Qualitatively, this variant tends to converge to over-pruned policies that miss key evidence, suggesting that the recall-oriented warm-up in Stage I is crucial for exploring a diverse evidence space before learning to compress it. Taken together, the ablations show that both the redundancy-aware reward and the staged optimization scheme are necessary to realize the full benefits of Context-Picker.
Method Judge Acc (%) $\Delta$ vs. full
Context-Picker (full) 70.6
   w/o rationale 64.1 6.5-6.5
   w/o redundancy 66.0 4.6-4.6
   w/o Stage 1 56.5 14.1-14.1
Table 2: Ablation study of Context-Picker on LoCoMo. $\Delta$ denotes absolute drops in Judge Acc (percentage points) compared to the full model.
5 Related Works
5.1 Adaptive Retrieval and Context Optimization
Standard RAG systems typically retrieve a fixed top-$K$ set of passages using sparse or dense retrievers (robertson2009probabilistic; karpukhin-etal-2020-dense; lewis2021retrievalaugmentedgenerationknowledgeintensivenlp; izacard2022atlasfewshotlearningretrieval), often combined with cross-encoder or sequence-to-sequence rerankers such as monoT5 to improve ordering quality (nogueira2020documentrankingpretrainedsequencetosequence; sun2024chatgptgoodsearchinvestigating; drozdov2023paradepassagerankingusing; chen2025scirerankbenchbenchmarkingrerankersscientific). While this “retrieve-then-rerank” architecture substantially improves recall and ranking, it still relies on a static $K$ for downstream generation. As a result, complex multi-hop questions may suffer from missing evidence when $K$ is small, whereas simple factoid queries incur unnecessary noise and cost when $K$ is large, exacerbating long-context issues such as distractor accumulation and the “lost-in-the-middle” effect (liu2023lostmiddlelanguagemodels; jin2024longcontextllmsmeetrag). Recent analyses of long-context RAG pipelines further show that simply increasing the number of retrieved passages often yields higher recall but only marginal or even negative gains in answer accuracy (jiang2024longragenhancingretrievalaugmentedgeneration; jin2024longcontextllmsmeetrag).
To overcome the rigidity of fixed-size retrieval, a series of works have explored more adaptive strategies. Self-RAG (asai2023selfraglearningretrievegenerate) trains a single LM augmented with reflection tokens to decide, segment by segment, when to retrieve, when to critique evidence, and when to continue generation. FLARE (jiang2023activeretrievalaugmentedgeneration) performs active retrieval by monitoring low-confidence tokens and issuing retrieval queries only when the model anticipates future uncertainty. Adaptive-RAG (jeong2024adaptiveraglearningadaptretrievalaugmented) introduces a query-complexity classifier that routes questions to no-retrieval, single-step, or iterative RAG pipelines, and Adaptive-$k$ chooses the number of selected passages from the similarity-score distribution of candidates without additional tuning or iteration (taguchi2025efficientcontextselectionlongcontext). These methods show that adjusting when and how much to retrieve can improve overall QA performance, but they typically require multiple rounds of retrieval and generation or rely on hand-crafted decision rules rather than an explicitly learned selection policy under a token budget.
Another line of work targets the context side of the pipeline via compression. LLMLingua (jiang2023llmlinguacompressingpromptsaccelerated) uses a smaller model to score and remove non-essential tokens inside prompts, yielding substantial speedups while preserving task performance. RECOMP (xu2023recompimprovingretrievalaugmentedlms) compresses retrieved documents into concise textual summaries before feeding them to the generator, reducing both prompt length and the burden on the LM to locate relevant information. These approaches operate primarily at the token or sentence level and focus on shrinking a given context, without explicitly reasoning about which subset of passages is minimally sufficient for answering the query.
Complementary to these advances, several works study context selection from a scoring perspective. Query-aware and list-aware rerankers use LLMs or specialized models to assign relevance scores to passages individually or jointly, sometimes with list-wise prompting that considers redundancy and coverage (sun2024chatgptgoodsearchinvestigating; chen2025scirerankbenchbenchmarkingrerankersscientific). Generator-aware metrics evaluate how well candidate contexts align with the generator’s internal knowledge or its preferences via reward models trained from LLM feedback (wang2024learningretrieveincontextexamples). Influence-guided selection goes one step further and defines a leave-one-out style Contextual Influence Value that measures performance degradation when removing each passage (deng2025influenceguidedcontextselection). However, most of these methods still operate at the level of per-passage utilities plus thresholding, and are not designed to directly optimize for a minimal sufficient subset under a strict input budget.
5.2 Reinforcement Learning for Evidence Selection
Reinforcement learning (RL) has been widely used to align retrieval-augmented systems with downstream tasks. Early work applied RL to optimize query reformulation, where an agent learns to rewrite user queries to better exploit a fixed retrieval module (zhu2025convsearchr1enhancingqueryreformulation), or to train retrievers end-to-end with task feedback, as in REALM-style frameworks that update both the encoder and retriever to maximize QA reward (guu2020realmretrievalaugmentedlanguagemodel). These approaches improve the quality of retrieved candidates, but still leave the final context selection to static Top-$K$ heuristics or simple truncation.
More recent approaches bring RL or RL-inspired feedback closer to the evidence and memory selection step itself. DynamicRAG (sun2025dynamicragleveragingoutputslarge) models the reranker as an agent over document sequences and trains it with a combination of supervised fine-tuning and RL, using LLM-judged answer quality as reward to adjust both the order and the number of retrieved documents. Memory-R1 and related memory agents frame long-term memory management and retrieval decisions as RL problems, training policies to decide what to store, update, or retrieve in order to support downstream QA and dialogue over very long conversational histories (yan2025memoryr1enhancinglargelanguage; maharana2024evaluatinglongtermconversationalmemory). RL has also been applied to conversational query reformulation and retrieval alignment and to broader agentic RAG frameworks such as RAG-Gym and REX-RAG, which optimize multi-step retrieval and reasoning trajectories with policy-gradient–style updates (zhu2025convsearchr1enhancingqueryreformulation; xiong2025raggymsystematicoptimizationlanguage; jiang2025rexragreasoningexplorationpolicy). Influence-guided context selection (deng2025influenceguidedcontextselection) employs a generator–judge loop to estimate each passage’s marginal influence via leave-one-out utilities and then trains a surrogate selector to approximate these influence scores.
While these RL-style or RL-adjacent approaches introduce valuable task-aligned signals, they still face two key challenges for context selection: (i) rewards are largely trajectory-level and sparse, as the agent receives only a scalar signal after producing a full list or trajectory, making credit assignment to individual passages and redundancy penalties difficult; and (ii) policies are typically optimized to improve list-wise ranking quality, to include all positively-scored contexts, or to manage memory operations, rather than to identify a minimal evidence subset that preserves answerability under a fixed input budget. In contrast, Context-Picker is trained on offline-mined minimal sufficient evidence sets and uses a two-stage, redundancy-aware GRPO objective to explicitly trade off coverage and compactness at the passage subset level.
@@ 6 Conclusion
In this work, we presented Context-Picker, a reasoning-aware framework that learns a variable-length evidence set under a token budget. Context-Picker combines (i) an offline evidence mining pipeline that distills greedily minimal sufficient evidence sets via a generator–judge loop with Leave-One-Out (LOO) pruning, providing dense and task-aligned supervision; and (ii) a two-stage reinforcement learning schedule optimized with GRPO, where Stage I (recall-oriented) emphasizes coverage of reasoning chains with a relaxed redundancy margin, and Stage II (precision-oriented) tightens redundancy penalties to prune distractors and distill compact support sets. The picker further outputs a rubric-guided rationale together with selected passage IDs, enabling structured, end-to-end consistent evidence selection.
Experiments on five long-context and multi-hop QA benchmarks demonstrate that Context-Picker outperforms strong RAG baselines under comparable evidence budgets, achieving higher LLM-as-judge accuracy while often with comparable or reduced context lengths. Ablation studies further verify that the coarse-to-fine schedule, redundancy-aware reward shaping, and rationale-guided output format each contribute substantially to the gains, and that removing Stage I leads to severe over-pruning and degraded performance. Future work may include extending Context-Picker to more open-ended generation tasks, exploring alternative reward signals beyond LLM-as-judge, and integrating the picker with token- or KV-level compression inside the generator to further reduce inference cost.
7 Acknowledgement
We gratefully acknowledge the support from the Distinguished Young Scholars Project funded by the Natural Science Foundation of Guangdong Province (No. 2025B1515020060), the Basic and Applied Basic Research Program of the Guangzhou Science and Technology Plan (No. 2025A04J7141).

### v2

@@ 4.2 Main Results
Performance comparison.
Table 1 presents the comparative results. Context-Picker (Stage 2) achieves the best performance on four out of five benchmarks, outperforming both standard RAG and advanced baselines. Specifically, on reasoning-heavy datasets like MuSiQue and HotpotQA, our method surpasses the strong reranking baseline RankRAG by margins of $5.7\%$ and $1.5\%$, respectively. This highlights that for multi-hop questions, simply reordering documents (RankRAG) is insufficient; explicitly pruning distractors to form a minimal evidence set is crucial for accurate reasoning. Compared to RL-based peers (DynamicRAG, Memory-R1), Context-Picker also demonstrates superior accuracy, particularly on the ultra-long LoCoMo benchmark.
Recall vs. Precision Trade-off.
On MultiFieldQA, RankRAG achieves the top score, slightly outperforming our Precision-Oriented Stage II. We attribute this to the dataset’s needle-in-a-haystack nature: many questions are closer to extractive factoid QA, where answerablity is dominated by whether a small number of critical snippets are retrieved, and the benefit of redundancy suppression is comparatively limited. Our Recall-Oriented Stage I reaches $0.873$, statistically tied with RankRAG, indicating that the coarse stage effectively maximizes evidence coverage. In contrast, Stage II enforces a tighter redundancy margin and explicitly optimizes evidence compression; while this is advantageous on reasoning-heavy benchmarks by removing distractors (e.g., MuSiQue), it can occasionally prune borderline but helpful cues, leading to a slight drop on MultiFieldQA. Overall, the two-stage design provides a practical controllable trade-off: Stage I prioritizes recall, while Stage II prioritizes compactness and reasoning density.
@@
Reinforcement learning on discrete text selection is often characterized by instability. However, thanks to our dense reward supervision mined via LOO and the GRPO algorithm, Context-Picker demonstrates robust training dynamics. As illustrated in Figure 3, the reward curves for both the Recall-Oriented Stage I (Above) and Precision-Oriented Stage II (Bottom) show steady convergence. The minimal gap between training and validation performance further validates the generalization capability of our offline evidence mining strategy.
@@ 4.3 Ablation Studies
In the full model, the picker outputs both a short natural-language rationale and a set of selected IDs. Removing the rationale branch (“w/o rationale”) leads to a $6.5$-point drop in Judge Acc. We hypothesize that requiring the model to verbalize why certain passages are selected acts as a structural regularizer: it encourages more stable reasoning over evidence interactions and reduces the tendency to over-select loosely related passages.
@@
Finally, we examine a variant trained only with the Stage II objective (“w/o Stage 1”), i.e., directly optimizing the refinement-oriented reward from scratch. This leads to a substantial degradation to $56.5\%$ Judge Acc, $14.1$ points below the full two-stage Context-Picker. Although the “w/o Stage 1” variant optimizes the same final reward function, it suffers from the cold-start problem inherent in combinatorial optimization. The Stage 2 reward imposes a strict penalty on redundancy. Without the high-recall initialization provided by Stage 1, the agent faces a sparse reward landscape: exploring larger subsets incurs immediate length penalties, while small random subsets rarely contain the complete evidence chain required for reasoning. Consequently, the policy gets trapped in a local optimum of selecting overly sparse contexts to minimize penalties, failing to answer the question.
These ablation studies show that our coarse-to-fine optimization schedule, the redundancy-aware reward shaping, along with the rationale generated by the policy, all contribute to our method.
5 Related Work
Adaptive RAG.
While standard RAG retrieves a fixed top-$K$ passages Lewis et al. (2021), recent works attempt to adapt context dynamically. Approaches like Self-RAG Asai et al. (2023) and FLARE Jiang et al. (2023) interleave retrieval with generation, but incur high inference latency. Adaptive-RAG Jeong et al. (2024) and Adaptive-$k$ Taguchi et al. (2025) select retrieval depth via classifiers or score thresholds, yet they rely on heuristics rather than optimizing for evidence sufficiency. Unlike rerankers that score passages individually Sun et al. (2024); Wang et al. (2024), Context-Picker optimizes the evidence subset directly.
RL for Retrieval.
RL has been applied to query reformulation Zhu et al. (2025) and retriever alignment Sun et al. (2025). However, methods like DynamicRAG typically optimize for ranking quality or answer correctness via sparse, trajectory-level rewards. Recent memory-agents Yan et al. (2025); Maharana et al. (2024) use RL for long-term memory management but often lack mechanisms to penalize redundancy explicitly. Our framework differs by using offline-mined “minimal sufficient sets” to provide dense supervision, enabling the policy to learn precise evidence compression.
@@ 6 Conclusion
We introduced Context-Picker, a reinforcement learning framework that shifts context retrieval from static ranking to dynamic, minimal sufficient subset selection. By leveraging offline-mined supervision and a two-stage coarse-to-fine training schedule, our model learns to balance high recall with strict compactness. Experiments across five benchmarks demonstrate that Context-Picker outperforms strong RAG baselines in answer accuracy. Future work will extend this paradigm to open-ended generation tasks and investigate integration with token-level compression techniques to further optimize inference costs.